# SupplementaryMethods Ultra-high dose rate 6 MeV electron irradiation generates stable [1-13C]alanine radicals suitable for medical imaging with dissolution Dynamic Nuclear Polarisation

### v1

@@ 2.3 In Vivo Application as a Metabolic Tracer
To test the optimised irradiated alanine preparation and DNP protocol, two proof of concept in vivo studies were conducted in rats. First, slice-selective magnetic resonance spectroscopy acquisitions were performed following separate tail-vein injections of [1-13C]alanine polarized with trityl radicals as described previously[44] or endogenous radicals generated via 70 kGy electron-beam irradiation. Temporal spectroscopic analysis demonstrated higher initial signal magnitude from the irradiated sample relative to the trityl-based preparation (Fig. 5a, b). Secondary resonances were observed flanking the primary [1-13C]alanine signal in both preparations. Spectral summation was performed to enhance the signal-to-noise ratio (SNR) and facilitate metabolite identification (Fig. 5c, d). The chemical shifts of these secondary resonances ( and ) corresponded precisely with the literature-reported values for [1-13C]lactate and [1-13C]pyruvate, respectively[45]. Subsequent metabolite quantification via the AMARES algorithm revealed comparable metabolic dynamics between both preparation methods (Fig. 5e, f) and spectral SNR was sufficient to quantify the downstream metabolites expected in vivo. No additional labelled peaks were observed; the products of the dissolution were identical to that observed with trityl radical.
Secondly, comparative 2D magnetic resonance spectroscopic imaging (MRSI) was performed in a separate animal using identical hyperpolarised [1-13C]alanine preparations. MRSI data were acquired, together with T2-weighted proton images acquired in the axial plane for anatomical localisation. Spatiotemporal analysis of the spectroscopic data demonstrated concordant biodistribution patterns between the trityl-based and irradiation-based preparations (Fig. 5g), with predominant accumulation in the renal parenchyma as alanine perfuses the kidneys, consistent with previous renal reports and enabling the technique to quantify redox potential,[44] and confirming equivalent in vivo behaviour between preparation methods.
@@ 3 Discussion
The clinical feasibility and utility of metabolic imaging with hyperpolarised MRI and dDNP has been consistently demonstrated since its inception 22 years ago [46], with the transformative ability to spatially map metabolism of huge interest in many disparate diseases including cancer,[1] heart disease,[47] and diabetes[48]. With more than 1000 human patients scanned and multi-centre trials soon to commence[49] it is a physical technique that has opened new avenues for research and the diagnosis and management of disease. However, for its integration within routine clinical workflows to be fully realised, significant improvements in its cost, accessibility and practicality are essential. Beyond the substantial capital expenditure for dDNP hardware (typically exceeding $2 million), institutions using dDNP face prohibitive ongoing operational costs. These include maintaining dedicated clean-room pharmacy facilities with specialised personnel for the sterile preparation of per-patient disposable fluid paths for dDNP, implementing expensive quality control measures, and adhering to stringent regulatory frameworks for pharmaceutical preparation. Each hyperpolarised examination currently requires the labour-intensive preparation of sterile fluid paths,[50] their filling with sterile stable-isotope labelled metabolites together with dissolution and neutralisation media (both manufactured according to sterile guidelines that vary regionally[49]) and a trityl radical that provides the unpaired electron the technique requires. The synthesis of trityl radicals is not trivial, typically requiring 13 separate organic synthetic steps, leading to a cost $10k/g.[6] Moreover, the required subsequent filtration steps to remove the radical add further complexity to the design of the fluid path. These are all individual processes that substantially increase per-patient costs and logistical complexity and restrict the availability of a technique that has the ability to directly measure central metabolic reactions key to all life. Furthermore, the limited shelf-life of the radical in prepared samples (approximately 24 hours at room temperature for trityl[51]) necessitates precise coordination between pharmacy operations and clinical scheduling, and prohibits the transport of dDNP samples nationally or internationally unless extraordinary steps are taken such as transporting samples cryogenically under defined (and varying) temperatures and magnetic fields.[52, 53, 12] As well as restraining human experiments, these large costs further constrain dDNP availability in non-human biological research: although of utility studying species such as snakes[45] and veterinary canine patients,[54] trityl radicals are sufficiently expensive as to be repurified from animal urine after use.[55]
@@
To address these fundamental constraints on widespread adoption, we explored ultra-high dose rate electron-beam irradiation for generating inherently biologically sterile samples with radicals in situ: a sample could be prepared in a central (worldwide) facility, rapidly sterilised, and shipped at room temperature around the world prior to use with a polariser in a hospital site. Furthermore, the ability to use powders or crystals of neutral salts, such as alanine, rather than glassing suspensions of trityl radicals and typically organic ketoacids, further dramatically ameliorates the technical requirements of the technique. This approach also opens new classes of molecules to dDNP, as radical formation occurs statistically throughout the sample via well-understood particle-physics interactions, rather than requiring glassed mixtures of disparate radicals (themselves being limited by exclusion from any crystals formed). This allows for effective polarisation of high concentrations of virtually any anhydrous solid-state molecule that would otherwise be challenging to study. All that would be required for dissolution is a neutral sterile dissolution buffer, and, unlike UV-irradiation generated radicals in acids, there is no need for precise control of temperature and magnetic field over many orders of magnitude.[12] Indeed, in this work, we prepared samples for DNP in one country (the UK) and studied them via dissolution-DNP in another (Denmark).
@@
As anticipated from extensive studies on NA alanine as a dosimeter, e-beam irradiation of NA alanine generated radicals in a dose-dependent manner[56], and a similar trend was observed from the 13C-enriched samples, saturating at around , well above the dose requirement of for medical sterilisation under ISO 11137-2.[57, 58] What is surprising, however, is that upon mixing with glycerol the system appears to become partially ordered, and the efficiency of DNP increases substantially. This also generates a clinically useful probe as alanine and glycerol are not miscible in the volume ratios considered here and stable alanine radicals remain present for at least 16 weeks, stored anhydrously in an inexpensive plastic container with silica desiccant gel at room temperature. Furthermore, changes at low temperature with this system, indicating a fundamental change in the radical electronic environment. We observed no unexpected radiation chemistry byproducts (likely because they spontaneously recombine upon the addition of hot water in the dissolution process, wherein the mixture is melted), and both glycerol and alanine are endogenous biomolecules – which are safe in vivo. This permitted the quantification in the rat kidney of downstream alanine metabolites, directly analogously to the reported biomedical uses of [1-13C]alanine hyperpolarised with trityl radicals.[59, 60, 44, 61, 62]
This polarisation enhancement is difficult to quantitatively understand. We have discovered a unique DNP mechanism that is not explained by conventional models yet produces clinically relevant levels of polarisation in an extensively studied biomolecule. The alteration of the frequency sweep profile of nuclear enhancement with radiation dose we hypothesise may be due to some degree of cooperativity between electrons. The temperature-dependent electronic -shift we observed in the EPR data raises intriguing questions about potential magnetic ordering in our system. Indeed, spontaneous spatial ordering in neutral alanine crystallites has previously been reported to occur in strong magnetic fields,[63] arising out of magnetic anisotropy of the L-alanine unit cell and its comparatively large (in magnitude) ;[64, 65, 66] and zirconium-nitrate doped -alanine single crystals form magnetic order and have been used as nonlinear optical devices.[42]
@@
While traditional paramagnetic systems maintain independent spin behaviour even at low temperatures, our partially ordered alanine crystallites dispersed in glycerol could exhibit more complex magnetic phenomena as radical centres become close enough to develop short-range magnetic correlations at low temperatures and high radiation doses within individual crystallites. The glycerol matrix in this scenario could function analogously to domain boundaries in ferromagnetic materials. By physically separating the alanine crystallites, glycerol creates discontinuities in the exchange pathways, limiting the correlation length of any magnetic ordering, but permitting thermodynamically driven spin diffusion through weak coupling provided by proton/carbon spins in the matrix. This is conceptually similar to the role of Bloch walls in ferromagnetic materials, though operating at the nano- or microscale between distinct crystallite domains rather than within a continuous bulk material. This model could potentially explain several of our observations:
@@
The levels of nuclear enhancement we achieved are several orders of magnitude larger than those reported for ionising-radiation radical induced DNP previously by -rays by Giannoulis,[10] and are comparable with the optimal trityl, narrow line-width radicals used clinically. For a non-irradiated alanine sample, the polarisation level at the point of dissolution was 19.2% and the nuclear T1 was . For reference, a previous study where alanine was also hyperpolarised with the trityl radical yielded a liquid state polarisation of 12.6% at the point of dissolution and an in vitro T1 of at [60]. Our irradiated samples produced comparable results; polarisation levels of 17.7% at the point of dissolution were achieved, while the extracted nuclear T1 relaxation times were at . The nuclear T1 relaxation time was prolonged due to the fact that the irradiation-generated radicals quench upon dissolution. Future work could explore the direct deuteration of the 13C-enriched alanine to potentially enhance polarisation levels as previously observed for other molecules hyperpolarised by dDNP[67]. Crucially, the pH of the irradiated alanine samples (approximately 7) also made them suitable for in vivo trial with dissolution in water; we note that neutral salts are readily crystallised and substances such as sodium pyruvate likely will be readily amenable to this technique
@@
Beyond the physical novelty of our approach, our in vivo experiments demonstrate significant translational potential. Electron-irradiated 13C-alanine not only matched but in some aspects exceeded the performance of conventional preparations using exogenous radicals. It is worth considering how 13C-alanine could compare to 13C-pyruvate, the most commonly used metabolic probe with dDNP and undergoing clinical trials. Whilst pyruvate polarises very well with dDNP and is rapidly metabolised to lactate, alanine, and bicarbonate, it has been noted that the pyruvate/lactate ratio is not reflective of intracellular redox state following HP pyruvate infusion[68]. Upon switching the injectable substrate from 13C-pyruvate to 13C-alanine, Hu et al. recorded more than a ten-fold increase in the 13C-lactate/13C-pyruvate ratio[60]. The difference was suggested to be due to the uptake of 13C-pyruvate being limited by perfusion and downstream flux through monocarboxylate transporters during the acquisition window. An incomplete uptake of 13C-pyruvate would mean contributions from an extracellular 13C-pyruvate pool in the measured 13C-pyruvate signal. Furthermore, any extracellular 13C-pyruvate present in plasma can be converted into 13C-lactate by red blood cells that contain lactate dehydrogenase, leading to an “extracellular” 13C-lactate pool[59]. Signals arising from intracellular and extracellular pools cannot be easily distinguished in hyperpolarised MR, leading to the proposed use of 13C-alanine for determining the 13C-lactate/13C-pyruvate ratio. For 13C-alanine to be converted into 13C-pyruvate, and subsequently 13C-lactate, it must first be exposed to the intracellular enzyme alanine transaminase. Thus, the 13C-pyruvate signal detected following the injection of 13C-alanine would be intracellular in origin. Moreover, after the metabolism of alanine is observed and quantified, the cellular redox potential can be easily determined [59] from the known relation where .
@@
Our electron irradiation approach uniquely combines this biochemical advantage with the practical benefits of in situ radical generation and spontaneous quenching, providing a powerful new tool for quantifying intracellular redox potential. It is also the first to explicitly consider partially ordered solids as substrates for biomedical dDNP rather than amorphous glasses that are commonly used with the technique (and, indeed, often listed as a requirement for its success; [69]) While this technique could be extended to other substrates, 13C-alanine represents an ideal initial candidate due to its metabolic significance, detailed quantum-mechanical study, and demonstrated biochemical and clinical relevance. This approach may reduce the technological, regulatory, and economic barriers that have constrained hyperpolarised MRI primarily to specialised research centres, potentially expanding access to this powerful diagnostic capability.
@@ 3.1 Limitations
We should note that several limitations should be acknowledged in this work: notably, our mechanistic understanding of the DNP process in this system remains incomplete and in need of extension via theoretical modelling. The unusual frequency sweep profiles of this highly efficient polarisation process hint at cooperativity and physics beyond conventional DNP models, but a comprehensive theoretical framework that quantitatively predicts these behaviors is lacking. Additional spectroscopic studies with variable temperature, field strength, and radical concentration would help elucidate the underlying mechanisms. Additionally, while we demonstrated stable radical formation in alanine, the applicability of this approach to other metabolically relevant molecules requires further investigation. Different molecular structures may respond differently to electron irradiation, potentially generating distinct radical species with varying stability and DNP efficiency. The specific arrangement of alanine in a crystalline lattice likely contributes to the observed radical stability and its partial ordering in glycerol (analogous to a nemantic liquid crystal), and molecules with different crystal structures, or those that do not readily crystallise, may present challenges. Furthermore, the comparatively high radiation doses considered may embrittle plastics used in the construction of fluid paths, but this phenomenon is well-characterised, and many medical plastics are routinely sterilised with comparable doses.[70, 71]
@@
Finally, the highest nuclear polarisation levels achieved with our approach () remain lower than those reported for optimised pyruvic acid preparations with trityl radicals (). While sufficient for in vivo imaging, higher polarisation would improve SNR and potentially allow detection of lower-concentration metabolites. Further optimisation of irradiation parameters, alanine crystallite size, and DNP conditions might narrow this gap. Likewise, owing in part to this limitation, our in vivo experiments were limited to proof-of-concept demonstrations in a small number of healthy animals. The performance of irradiated alanine in disease models, particularly its sensitivity to pathological alterations in redox potential, requires validation in relevant preclinical models before clinical translation.
4 Methods

### v2

@@ 2.3 In Vivo Application as a Metabolic Tracer
To test the optimised irradiated alanine preparation and DNP protocol, two proof of concept in vivo studies were conducted in rats. First, slice-selective magnetic resonance spectroscopy acquisitions were performed following separate tail-vein injections of [1-13C]alanine polarised with trityl radicals as described previously (?) or endogenous radicals generated via 70 kGy electron-beam irradiation. Temporal spectroscopic analysis demonstrated higher initial signal magnitude from the irradiated sample relative to the trityl-based preparation (Fig. 5A, B). Secondary resonances were observed flanking the primary [1-13C]alanine signal in both preparations. Spectral summation was performed to enhance the signal-to-noise ratio (SNR) and facilitate metabolite identification (Fig. 5C, D). The chemical shifts of these secondary resonances (183 ppm and 171 ppm) corresponded precisely with the literature-reported values for [1-13C]lactate and [1-13C]pyruvate, respectively (?, ?, ?, ?). Subsequent metabolite quantification via the AMARES algorithm revealed comparable metabolic dynamics between both preparation methods (Fig. 5E, F) and spectral SNR was sufficient to quantify the downstream metabolites expected in vivo, although the comparatively long TR of 3 s precluded a detailed study of downstream kinetics whilst preserving magnetisation. No additional labelled peaks were observed; the products of the dissolution were identical to that observed with trityl radical (c.f. SI §6.14). The dissolution was ‘clean’, meaning that the entire volume of the sample was dissolved by the 5 ml D2O used as a dissolution buffer with no remaining partially dissolved sample remaining, and a final concentration of approximately 100 mM dissolved alanine was obtained post-dissolution, and subsequently 1 ml was manually injected over 5 s.
[Figure 5 approximately here]
Secondly, comparative 2D magnetic resonance spectroscopic imaging (MRSI) was performed in a separate animal using identical hyperpolarised [1-13C]alanine preparations. MRSI data were acquired, together with -weighted proton images acquired in the axial plane for anatomical localisation. Spatiotemporal analysis of the spectroscopic data demonstrated concordant biodistribution patterns between the trityl-based and irradiation-based preparations (Fig. 5G), with predominant accumulation in the renal parenchyma as alanine perfuses the kidneys, consistent with previous renal reports and enabling the technique to quantify redox potential, (?) and confirming equivalent in vivo behaviour between preparation methods.
@@ 3 Discussion
The clinical feasibility and utility of metabolic imaging with hyperpolarised MRI and dDNP has been consistently demonstrated since its inception 22 years ago (?), with the transformative ability to spatially map metabolism of huge interest in many disparate diseases including cancer, (?) heart disease, (?) and diabetes (?). With more than 1000 human patients scanned and multi-centre trials soon to commence (?) it is a physical technique that has opened new avenues for research and the diagnosis and management of disease. However, for its integration within routine clinical workflows to be fully realised, significant improvements in its cost, accessibility and practicality are essential. Beyond the substantial capital expenditure for dDNP hardware (typically exceeding $2 million), institutions using dDNP face prohibitive ongoing operational costs. These include maintaining dedicated clean-room pharmacy facilities with specialised personnel for the sterile preparation of per-patient disposable fluid paths for dDNP, implementing expensive quality control measures, and adhering to stringent regulatory frameworks for pharmaceutical preparation. Each hyperpolarised examination currently requires the labour-intensive preparation of sterile fluid paths, (?) their filling with sterile stable-isotope labelled metabolites together with dissolution and neutralisation media (both manufactured according to sterile guidelines that vary regionally (?)) and a trityl radical that provides the unpaired electron the technique requires. The synthesis of trityl radicals is not trivial, typically requiring 13 separate organic synthetic steps, leading to a cost $10k/g. (?) Moreover, the required subsequent filtration steps to remove the radical add further complexity to the design of the fluid path. The trityl radical used in human studies, AH111501, also known as EPA, is the methylated form of OX063, and is specifically engineered to precipitate out in acidic aqueous solution for filtration within the sterile fluid path. Indeed, to satisfy regulatory considerations it is currently a release criterion for all sites performing human hyperpolarised experiments that the radical concentration is below an agreed number, typically less than 3 mM in the dissolved liquid. (?) These are all individual processes that substantially increase per-patient costs and logistical complexity and restrict the availability of a technique that has the ability to directly measure central metabolic reactions key to all life.
Furthermore, the limited shelf-life of the radical in prepared samples (approximately 24 hours at room temperature for trityl (?)) necessitates precise coordination between pharmacy operations and clinical scheduling, (?) and prohibits the transport of dDNP samples nationally or internationally unless extraordinary steps are taken such as transporting samples cryogenically under defined (and varying) temperatures and magnetic fields. (?, ?, ?) As well as restraining human experiments, these large costs further constrain dDNP availability in non-human biological research: although of utility studying species such as snakes (?) and veterinary canine patients, (?) trityl radicals are sufficiently expensive as to be repurified from animal urine after use. (?)
@@
To address these fundamental constraints on widespread adoption, we explored ultra-high dose rate electron-beam irradiation for generating inherently biologically sterile samples with radicals in situ: a sample could be prepared in a central (worldwide) facility, rapidly sterilised, and shipped at room temperature around the world prior to use with a polariser in a hospital site. Furthermore, the ability to use powders or crystals of neutral salts, such as alanine, rather than glassing suspensions of trityl radicals and typically organic ketoacids, further dramatically ameliorates the technical requirements of the technique, as a dissolution with sterile water is easier to obtain than with a carefully measured equimolar quantity of neutralising (basic) sterile buffer. This approach also opens new classes of molecules to dDNP, as radical formation occurs statistically throughout the sample via well-understood particle-physics interactions, rather than requiring glassed mixtures of disparate radicals (themselves being limited by exclusion from any crystals formed). This allows for effective polarisation of high concentrations of virtually any anhydrous solid-state molecule that would otherwise be challenging to study. All that would be required for dissolution is a neutral sterile dissolution buffer, and, unlike UV-irradiation generated radicals in acids, there is no need for precise control of temperature and magnetic field over many orders of magnitude. (?) Indeed, in this work, we prepared samples for DNP in one country (the UK) and studied them via dissolution-DNP in another (Denmark).
@@
As anticipated from extensive studies on NA alanine as a dosimeter, e-beam irradiation of NA alanine generated radicals in a dose-dependent manner (?), and a similar trend was observed from the 13C-enriched samples, saturating at around 70 kGy, well above the dose requirement of 25 kGy for medical sterilisation under ISO 11137-2. (?, ?) What is surprising, however, is that upon mixing with glycerol the system appears to become partially ordered, and the efficiency of DNP increases substantially. This also generates a clinically useful probe as alanine and glycerol are not miscible in the volume ratios considered here and stable alanine radicals remain present for at least 16 weeks, stored anhydrously in an inexpensive plastic container with silica desiccant gel at room temperature. Furthermore, changes at low temperature with this system, indicating a fundamental change in the radical electronic environment. We observed no unexpected radiation chemistry byproducts (likely because they spontaneously recombine upon the addition of hot water in the dissolution process, wherein the mixture is melted), and both glycerol and alanine are endogenous biomolecules – which are safe in vivo. This permitted the quantification in the rat kidney of downstream alanine metabolites, directly analogously to the reported biomedical uses of [1-13C]alanine hyperpolarised with trityl radicals. (?, ?, ?, ?, ?)
@@
This polarisation enhancement is difficult to quantitatively understand. We have discovered a unique DNP mechanism that is not explained by conventional models yet produces clinically relevant levels of polarisation in an extensively studied biomolecule. The alteration of the frequency sweep profile of nuclear enhancement with radiation dose we hypothesise may be due to some degree of cooperativity (interactions) between electrons when the radical concentration becomes sufficiently high that the picture of an isolated, well-defined radical species as an independent molecule breaks down. The temperature-dependent electronic -shift we observed in the EPR data is consistent with magnetic ordering in our system, or at the very least some degree of electronic reconfiguration over molecular length scales (c.f. Fig. S6), leading to a superposition of effective radical species that may interconvert and interact with neighbouring molecules (alanine has four in the unit cell), including the primary electronic radical disfavoured at high temperature. Indeed, spontaneous spatial ordering in neutral alanine crystallites has previously been reported to occur in strong magnetic fields, (?) arising out of magnetic anisotropy of the L-alanine unit cell and its comparatively large (in magnitude) ; (?, ?, ?) and zirconium-nitrate doped -alanine single crystals form magnetic order and have been used as nonlinear optical devices. (?) While traditional paramagnetic systems maintain independent spin behaviour and molecular character even at low temperatures, our partially ordered alanine crystallites dispersed in glycerol could exhibit more complex magnetic phenomena as radical centres become close enough to develop short-range magnetic correlations at low temperatures and high radiation doses within individual crystallites. The glycerol matrix in this scenario could function analogously to domain boundaries in ferromagnetic materials. By physically separating the alanine crystallites, glycerol creates discontinuities in the exchange pathways, limiting the correlation length of any magnetic ordering, but permitting thermodynamically driven spin diffusion through weak coupling provided by proton/carbon spins in the matrix. This is conceptually similar to the role of Bloch walls in ferromagnetic materials, though operating at the nano- or microscale between distinct crystallite domains rather than within a continuous bulk material. This model could potentially explain several of our observations:
@@
The levels of nuclear enhancement we achieved ( or at 1.4T) compare favourably to the to 800-fold enhancements reported at 11.7T for ionising-radiation radical induced DNP previously by Co-60 -rays by Giannoulis, (?) and are comparable with the optimal trityl, narrow line-width radicals used clinically. For a non-irradiated alanine sample prepared with trityl radical, the polarisation level at the point of dissolution was 19.2% and the nuclear T1 was 40.8 s at 1.4T. In contrast to the work by Giannoulis, we did not observe data consistent with a simple model by the solid-effect, and obtained qualitatively different frequency-sweep curves. This may be due to differences in the radiation transport physics of ultra-high dose rate 6 MeV electrons vs 1 MeV -rays, which will deposit more charge into the crystallites, as well as dose rate effects altering the environment around the radical centre, by, for example, introducing a degree of unresolved crystallographic strain. (?, ?, ?) We also note that Giannoulis’ alanine samples were flame-sealed and irradiated in vacuo (our samples were in air), and also polarised as dry powders. Finally, for reference, a previous study where alanine was also hyperpolarised with the trityl radical yielded a liquid state polarisation of 12.6% at the point of dissolution and an in vitro T1 of 41.5 s at 3 T (?). Our irradiated sample produced comparable results; polarisation levels of 17.7% at the point of dissolution were achieved, while the extracted nuclear T1 relaxation times were 77.3 s at 1.4 T. The nuclear T1 relaxation time was prolonged due to the fact that the irradiation-generated radicals quench upon dissolution. Future work could explore the direct deuteration of the 13C-enriched alanine to potentially enhance polarisation levels as previously observed for other molecules hyperpolarised by dDNP (?). Crucially, the pH of the irradiated alanine samples (approximately 7) also made them suitable for in vivo trial with dissolution in water; we note that neutral salts are readily crystallised and substances such as sodium pyruvate likely will be readily amenable to this technique.
@@
Beyond the physical novelty of our approach, our in vivo experiments demonstrate significant translational potential. Electron-irradiated 13C-alanine not only matched but in some aspects exceeded the performance of conventional preparations using exogenous radicals. It is worth considering how 13C-alanine could compare to 13C-pyruvate, the most commonly used metabolic probe with dDNP and undergoing clinical trials. Whilst pyruvate polarises very well with dDNP and is rapidly metabolised to lactate, alanine, and bicarbonate, it has been noted that the pyruvate/lactate ratio is not reflective of intracellular redox state following HP pyruvate infusion (?). Upon switching the injectable substrate from 13C-pyruvate to 13C-alanine, Hu et al. recorded more than a ten-fold increase in the 13C-lactate/13C-pyruvate ratio (?). The difference was suggested to be due to the uptake of 13C-pyruvate being limited by perfusion and downstream flux through monocarboxylate transporters during the acquisition window. An incomplete uptake of 13C-pyruvate would mean contributions from an extracellular 13C-pyruvate pool in the measured 13C-pyruvate signal. Furthermore, any extracellular 13C-pyruvate present in plasma can be converted into 13C-lactate by red blood cells that contain lactate dehydrogenase, leading to an “extracellular” 13C-lactate pool (?). Signals arising from intracellular and extracellular pools cannot be easily distinguished in hyperpolarised MR, leading to the proposed use of 13C-alanine for determining the 13C-lactate/13C-pyruvate ratio. For 13C-alanine to be converted into 13C-pyruvate, and subsequently 13C-lactate, it must first be exposed to the intracellular enzyme alanine transaminase. Thus, the 13C-pyruvate signal detected following the injection of 13C-alanine would be intracellular in origin. Moreover, after the metabolism of alanine is observed and quantified, the cellular redox potential can be easily determined (?) from the known relation where M. If a measure or estimate of intracellular pH is obtained – for example, from 31P-MRS (?) or through the use of hyperpolarised pyruvate (?) – this would provide a direct mechanism by which the NAD/NADH ratio could be non-invasively determined, as has been reported previously for trityl-hyperpolarised alanine. (?)
@@
Our electron irradiation approach uniquely combines this biochemical advantage with the practical benefits of in situ radical generation and spontaneous quenching, providing a powerful new tool for quantifying intracellular redox potential. It is also the first to explicitly consider partially ordered solids as substrates for biomedical dDNP rather than amorphous glasses that are commonly used with the technique (and, indeed, often listed as a requirement for its success; (?)) While this technique could be extended to other substrates, 13C-alanine represents an ideal initial candidate due to its metabolic significance, detailed quantum-mechanical study, and demonstrated biochemical and clinical relevance. This approach may reduce the technological, regulatory, and economic barriers that have constrained hyperpolarised MRI primarily to specialised research centres, potentially expanding access to this powerful diagnostic capability.
@@ 3.1 Limitations
We should note that several limitations should be acknowledged in this work: notably, our mechanistic understanding of the DNP process in this system remains incomplete and in need of extension via theoretical modelling. The unusual frequency sweep profiles of this highly efficient polarisation process hint at cooperativity and physics beyond conventional DNP models, but a comprehensive theoretical framework that quantitatively predicts these behaviors is lacking. Additional spectroscopic studies with variable temperature, field strength, and radical concentration would help elucidate the underlying mechanisms.
Additionally, while we demonstrated stable radical formation in alanine, the applicability of this approach to other metabolically relevant molecules requires further investigation. Different molecular structures may respond differently to electron irradiation, potentially generating distinct radical species with varying stability and DNP efficiency. The specific arrangement of alanine in a crystalline lattice likely contributes to the observed radical stability and its partial ordering in glycerol (analogous to a nemantic liquid crystal), and molecules with different crystal structures, or those that do not readily crystallise, may present challenges. Furthermore, the comparatively high radiation doses considered may embrittle plastics used in the construction of fluid paths, but this phenomenon is well-characterised, and many medical plastics are routinely sterilised with comparable doses. (?, ?) We chose to investigate alanine because it is exceptionally well studied from a radiation physics and EPR point of view, and is a reported biomolecular probe investigated for probing central metabolism. Other, perhaps more commonly employed biomolecules may well be amenable to this technique, where significant further study across all aspects of this highly interdisciplinary science remain yet to be elucidated. We can propose three fundamentally distinct approaches to explore this in future work: (1) by directly irradiating other molecules; (2) by irradiating alanine powder containing microcrystallites (such as that here) and dispersing alanine radicals in the solid state with other molecules of interest; or (3) by utilising co-crystallised (e.g. natural abundance) alanine and an isotopically labelled molecule of interest, under the assumption that radical species formed act as dilute spin impurities in the resulting system. We have yet to explore any of these other exciting avenues for the technique.
@@
Finally, the highest nuclear polarisation levels achieved with our approach () remain lower than those reported for optimised pyruvic acid preparations with trityl radicals (). While sufficient for in vivo imaging, higher polarisation would improve SNR and potentially allow detection of lower-concentration metabolites. Further optimisation of irradiation parameters, alanine crystallite size, and DNP conditions might narrow this gap. Likewise, owing in part to this limitation, our in vivo experiments were limited to proof-of-concept demonstrations in two healthy animals, precluding a detailed characterisation of the technical vs biological variability of the technique beyond providing proof-of-principle. The performance of irradiated alanine in disease models, particularly its sensitivity to pathological alterations in redox potential, requires validation in relevant preclinical models before clinical translation. We also note that the applicability of this technique may well be higher in other diseases and organs where alanine metabolism is believed to be more avid; for example, in the liver where the lactate/pyruvate ratio is closer to 10 following the infusion of alanine (rising to 20 after an ethanol challenge). (?)
Beyond scientific limitations, there is much that remains to be proven for the clinical adoption of this technique. At present, most European sites conducting human hyperpolarised experiments use electron-beam irradiation to sterilise (25-45 kGy) empty ‘sterile fluid paths’ made of polyethyletherketone (PEEK) and polyphenylsulfone (PPSU) plastics, which are filled under sterile conditions by a pharmacy team with pyruvic acid/trityl radical, snap frozen in liquid nitrogen, purged with helium under pressure, laser welded, and transported to the polariser for ultimate human use. (?) We envisage that filling a fluid path with reagents, irradiating the solid isotopically labelled compound to a higher level and the fluid path overall to a level for sterilisation would be a feasible method to generate transportable, sealed single-use devices in a centralised “worldwide” factory with an electron accelerator. However, we have not explicitly demonstrated this, nor have we shown that PEEK or PPSU walled tubes could withstand the high radiation doses required. We likewise have not undertaken the detailed preclinical toxicity analyses to show that no byproducts or adjuncts would be formed by high energy electron irradiation of PEEK or PPSU plastics containing solid isotopically labelled crystallite alanine, all of which is future work and necessitate completion in a validated, accredited laboratory prior to human use.
4 Materials and Methods